# Token Spammers, Rug Pulls, and SniperBots: An Analysis of the Ecosystem of Tokens in Ethereum and the Binance Smart Chain (BNB)

### v2

@@ Abstract
In this work, we perform a longitudinal analysis of the BNB Smart Chain and Ethereum blockchain from their inception to March 2022. We study the ecosystem of the tokens and liquidity pools, highlighting analogies and differences between the two blockchains. We estimate the lifetime of the tokens, discovering that about 60% of them are active for less than one day. Moreover, we find that 1% of addresses create an anomalous number of tokens (between 20% and 25%). We present an exit scam fraud and quantify its prevalence on both blockchains. We find that token spammers use short lifetime tokens as disposable tokens to perpetrate these frauds serially. Finally, we present a new kind of trader bot involved in these activities, and we detect their presence and quantify their activity in the exit scam operations.
@@ 1 Introduction
The cryptocurrency market is loosely regulated [33, 7]. Even if policymakers are moving towards building a safer environment for cryptocurrency investors [48], it is a complex task, and it needs time. Meanwhile, blockchain-related technologies evolve fast, and with the birth of the DeFi [57] investors begin to move from centralized exchanges (CEX) like Binance or FTX to decentralized exchanges (DEX). DEXes are distributed Applications (dApp) for trading that run on-chain powered by smart contracts. While regulating the standard cryptocurrencies market is not easy, ruling the on-chain trading platform is harder. Indeed, even if the web interface of a DEX can be shut down [5], its smart contracts are still reachable and working on the blockchain.
@@
DEX and DeFi dApp were born in the Ethereum blockchain, but DeFi services rapidly pop up on all the blockchains that support smart contracts. Although Ethereum always plays the role of the main character in the DeFi world, with over 68 billion USD locked in its smart contracts, the BNB smart chain or BSC (former Binance Smart Chain) proposes itself as a faster and cheaper alternative to Ethereum.
@@
Uniswap and PancakeSwap are arguably the two most popular DEXes on Ethereum and BSC. They rely on the Automated Market Maker (AMM) model to handle the trading system. At the basis of the AMM model, there is the concept of liquidity pools, a smart contract that handles two tokens (trading pair) that the user can swap. Unlike CEX, where the platform defines the trading pair, every user can create their pair on DEXes and let the other users use it. However, as we will see in the following, some users abuse this freedom to carry out frauds.
In this work, we perform a longitudinal investigation on tokens and liquidity pools that live in the Ethereum and BSC blockchains. We start by parsing over 3 billion transactions of both blockchains, searching for the largest possible number of tokens and liquidity pools. We identified more than 1.3 million tokens and 1 million liquidity pools. Then we reconstruct their lifetime, discovering that approximately 60% of the tokens have a lifetime shorter than 1 day. Focusing on who creates the token, we observe that a tiny fraction of addresses, just 1%, is responsible for creating more than 20% of the tokens. Given their overproduction of tokens, we will call these addresses token spammers. Surprisingly, we also find that the tokens with a very short life are actively traded on liquidity pools. Albeit this phenomenon is present on both blockchains, we notice that it is more widespread in BSC.
Diving into this subset of tokens, we observe that a large fraction of liquidity pools used to trade the 1-day tokens show the malicious pattern that we call exit scam. We analyze all the liquidity pools looking for exit scam patterns, and we find 266,340 potential frauds in BSC and 21,594 potential frauds in Ethereum. We estimate the cost of the operation and the gains of the scammers. Here, we see that the success rate of this operation is not very high (between 40% and 60%). However, given the simplicity and the very low cost of the operation, scammers can serially arrange the fraud and cover a series of unsuccessful operations with a single successful one.
@@
  • 1-day tokens and Token spammers: We estimate the lifetime of the tokens on both blockchains, discovering that about 60% of tokens last less than one day. In particular, a significant fraction of them lasts just 1 block. Analyzing who creates the tokens, we observe that just 1% of addresses create an abnormal number of tokens (about 20-25% of tokens of the blockchain).
  • Exit scam operations: We investigate the presence of the exit scam fraud pattern in 1-day tokens. We discover that in BSC, 81.2% of 1-day tokens listed on PancakeSwap have this pattern. Moreover, we notice that 75.1% of the token spammer in BSC carry out at least one operation using “disposable” tokens. We estimate the gains of the scammer, observing that even if the operation is very simple to arrange, given its cheap cost, it is profitable when performed serially.
  • Sniper bots: We find the presence of sniper bots, a particular kind of trader bot that observes the mempool of the blockchain to buy newly listed tokens in almost zero time from the moment they are tradable on a liquidity pool. To the best of our knowledge, we are the first to illustrate how this kind of trading bot works, detect their presence, and quantify their activity in the exit scam operations.
@@ 2.1 Ethereum
Ethereum [12] is a proof-of-work blockchain. Its native coin is the Ether (ETH), the second most popular cryptocurrency after Bitcoin (BTC) with a market cap of more than 210 billion US dollars. A key feature of Ethereum is smart contracts, pieces of code that can execute in a decentralized way on-chain, making Ethereum a programmable blockchain. Smart contracts enable the creation of decentralized applications (dApp) and the so-called Web3.0 [3]. Through smart contracts, it is possible to create new digital assets like (fungible) tokens and NFTs (not fungible tokens).
@@
The tokens. Tokens, like coins, are cryptocurrencies that can be exchanged or traded. The main difference is that coin is the native asset of the blockchain, whereas tokens are created on top of the blockchain, and their mechanisms are defined using smart contracts. In Ethereum, the ERC-20 [25] standard defines the main properties of tokens. ERC-20 was proposed in late 2015 to establish the standard interface for tokens. An ERC-20 compliant smart contract must implement a set of functions and events specified in the standard. These functions are reported in Table 1. Some of them are optional, in particular the name(), the symbol(), and the decimal() functions. In Ethereum, tokens and digital assets are held into accounts.
@@
Ethereum accounts. There are two kinds of accounts in Ethereum: Externally owned accounts (EOA) and contract accounts. EOAs consist of a pair of public and private keys generated with the Elliptic Curve Digital Signature Algorithm (ECDSA) [30]. An account is represented by its public address, a 42-character hexadecimal string obtained concatenating "0x" to the last 20 bytes of the Keccak-256 [22] hash of the public key. Generally, users interact with an account using applications called wallets. Example of wallets are MetaMask [40], TrustWallet [54] or MyEtherWallet [41]. A contract account, instead, is an account tied to a smart contract, and it is represented with an address of the same format as an EOA. A contract account is generated when a smart contract is deployed to the Ethereum blockchain. Both accounts can hold and send Ether. However, contract accounts can only send transactions in response to receiving a transaction.
@@
Transactions and fee. A transaction is an action that updates the whole Ethereum network. It can be used to move digital assets, deploy a smart contract, or invoke a smart contract. Executing a transaction has a cost, commonly called transaction fee. The fee is variable and depends on two main factors: The state of the network (if the network is heavily loaded, the fee is usually higher), and the complexity of the operation that the transaction triggers. For instance, moving Ether from an EAO to another is the cheapest kind of transaction, while interacting with a smart contract could be very expensive. For the sake of simplicity, we can say that the transaction fee is composed of two parts: the gas limit and the gas price. Gas refers to the unit that measures the computational effort required to execute specific operations. The gas limit represents the maximum amount of gas a user is willing to pay for the operation, and it has to be high enough to pay the computational effort; otherwise, the transaction will fail. Instead, the gas price is the amount of Gwei ($10^{-9}$ Ether) the user is willing to pay for each gas unit.
@@
Smart contract deployment. As we said, smart contracts are programs that run on the Ethereum blockchain. They are written in a high-level programming language (e.g., Solidity [20]) and compiled into bytecode that runs on the Ethereum Virtual Machine (EVM) [24]. A smart contract can be deployed by sending a contract creation transaction from an EOA to the zero address1110x0000000000000000000000000000000000000000. The transaction contains the bytecode of the smart contract. A smart contract can also create new smart contracts. In this case, the bytecode of the new smart contract has to be embedded in the bytecode of the smart contract that generates the new one. Since a smart contract can not start a transaction by itself but only in response to a transaction that triggers it, an EOA must trigger the generation of a new smart contract.
Events and logs. A smart contract has data associated with it, such as its Ether balance and the value of its variables. Transactions, calling the smart contract methods, can modify those values, hence the state of the smart contract itself. Knowing the internal state of a smart contract can be crucial, especially in cases where it serves as a backend to distributed applications (dApps). Ethereum provides Events and a Logs register to track the internal states of smart contracts. Each time an action changes the internal state of a smart contract, it can fire an Event that will notify the change. All the events are written on an Event log. Thanks to it, users and developers can easily track the state of the smart contracts in the blockchain.
@@
EVM and EVM compliant. Ethereum is a distributed state machine that changes its state at each new block accordingly to a predefined set of rules. The EVM is the entity that computes these changes in states. Specifications of the EVM are described in the Ethereum Yellowpaper [55]. There are several standard implementations of the EVM in different programming languages (e.g., Python, JavaScript, C++). Other than Ethereum, other blockchains rely on the EVM (to name a few: BNB Smart Chain [9], Avalanche [46], Fantom [26], Cronos [18]), and they use one of the standard EVM or a complete custom one. These blockchains are called EVM compliant. They run the same (or with minimal change) smart contract written in Ethereum, use the same convention for the address, and handle states the same way as Ethereum.
2.2 The BNB Smart Chain
@@
The BNB Smart Chain [9] (previously Binance Smart Chain) or BSC is a blockchain that was born in 2020 as a parallel to the Beacon Chain (previously Binance Chain), and together they form the BNB Chain. Its consensus is based on the PoSA [10] (Proof of Stake and Authority). While the Beacon chain handles the staking and the governance of the blockchain, the BSC manages the consensus layer and provides EVM compatibility. The coin of both chains is the BNB (Build and Build, previously Binance Coin)—the third coin by market cap with over 46 billion of capitalization. As Ether on Ethereum, the BNB coin fuels the transactions in the BNB chain. Given the EVM compatibility, creating tokens on the BSC is possible. However, in this case, tokens follow the BEP-20 standard instead of the ERC-20.
3 The Automated Market Maker - AMM
Cryptocurrencies can be traded on centralized exchanges (CEXs)—like Binance, Coinbase, FTX—or on decentralized exchanges (DEXs). The name CEX stems from the fact that users who trade a cryptocurrency for another have to transfer the cryptocurrency from their private wallet to a custodial wallet managed by a centralized entity, that is the exchange. Usually, on CEX, like on the traditional stock exchange, trades are performed following the Order Book Model. In this model, there is a system that matches the users’ trading orders, ensuring that each order is closed accordingly. In simple terms, if trader A wants to buy 1 Bitcoin at \$40,000, the system performs the trade only if it finds a trader B willing to sell 1 Bitcoin at the same price.
Unlike CEX, on DEX there is no intermediary. The user interacts with the smart contracts deployed on the blockchain, and the user’s cryptocurrencies leave their private wallet only when traded. Although some DEXs operate the Order Book Model (EtherDelta, Binance Dex, Waves Exchange, DyDx), the most popular follow the Automated Market Maker model. This model relies on a mathematical formula to fix the price of assets and on the concept of liquidity pools and liquidity providers. A liquidity pool is a smart contract that contains two or more cryptocurrencies that the user can swap one for the other. Instead, a liquidity provider is a user who invests in the liquidity pool, providing cryptocurrencies to the smart contract. When a liquidity provider injects liquidity into the liquidity pool, the smart contract mints LP-tokens and gives them to the liquidity provider. The LP-token represents the share of the liquidity pool owned by the investor. Conversely, when the liquidity provider desires to get back her cryptocurrencies, he transfers the LP-tokens to the smart contract. The latter burns the LP-tokens and provides the cryptocurrencies back to the investor. Usually, liquidity pools apply a trading fee to each swap operation and distribute a portion of the fees to the liquidity providers according to their LP-tokens.
3.1 Uniswap and its forks
Uniswap [1] is the first decentralized application (DEX) to use the AMM model successfully. According to DefiLlama [36], a popular DeFi statistics aggregator, Uniswap is the $5^{th}$ dApp by TVL (Total Value Locked, amount of money locked into smart contracts) with over 6 billion USD, while it is the $1^{th}$ among the AMMs. Uniswap was launched on Ethereum, but now it is also present on the Ethereum Layer 2 solutions Arbitrum and Optimism and on the Ethereum side chain Polygon Matic. Because of its popularity, its open-source smart contracts, and the copyleft license [51], more than 50 protocols were born on several blockchains by forking Uniswap smart contracts in the last years. Uniswap is on its third version, but all its forks belong to the second version since the third one is under a Business Source License [52]. For this reason, in this work, we focus on Uniswap V2 and its forks. One of the most popular fork of Uniswap is PancakeSwap, which lives in BSC, and it is the $1^{th}$ dApp by TVL on this blockchain with over 4 billion of USD locked in its smart contracts.
@@ 3.1 Uniswap and its forks
In Uniswap V2, each pool consists of a pair of ERC-20 tokens. We can think of the liquidity pool as divided into two parts, each containing a single token, and both have an equivalent value. Let a pool consist of $x$ token A and $y$ token B. At each swap, the pool preserves $xy$. When a user swaps $a$ token A for token B (the user adds token A to the pool and takes token B from the pool), $x$ increases by $a$ and $y$ decreases by $b$, where $b$ is computed so that $xy$ does not change. The rate $a/b$ of the exchange is determined by the ratio of $x$ and $y$ in the pool. Consequently, the swap operation changes the current exchange rate, the value of token A decreases while the value of token B increases, and the two parts maintain the same value.
Although Uniswap has some pools directly created by itself, anyone can leverage the protocol and create a new pool with a custom trading pair. This freedom allows the creator of the pool to establish the initial price of a new token, balancing the amount of liquidity added to the liquidity pool.
@@ 4 The Datasets
For our investigation, we build two different datasets: The Token Dataset contains all the ERC-20 (resp. BEP-20) tokens created, and the Liquidity Pool Dataset, contains data about liquidity pools. Each dataset has two versions, one with data from the Ethereum blockchain and the other from the BNB Smart Chain.
Given the large amount of data and the need to parse the entire blockchains multiple times, for performance reasons and to avoid overloading public nodes (e.g., nodes provided by Binance [8] and Infura [29]) or services (e.g., BSCscan or Etherscan), we host and run an Ethereum and a BNB Smart Chain node. Finally, to query the blockchains and process the data we use the Web3 [42] and the Ethereum-etl [39] Python libraries. Web3 is a library that allows interaction with a local or remote EVM-compliant node using HTTP, IPC, or WebSocket. Ethereum-etl allows extraction of information from EVM-compliant blockchain and exporting them into formats like CSV or JSON.
We consider the whole history of both blockchains from their inception to March 2022. For the Ethereum blockchain, we process all the blocks from block 0 (2015-07-30) to block 14340000 (2022-07-03). For the BSC blockchain from block 0 (2020-04-20) to block 15854000 (2022-03-07).

### v3

# Token Spammers, Rug Pulls, and Sniper Bots: An Analysis of the Ecosystem of Tokens in Ethereum and in the Binance Smart Chain (BNB)

@@ Abstract
In this work, we perform a longitudinal analysis of the BNB Smart Chain and Ethereum blockchain from their inception to March 2022. We study the ecosystem of the tokens and liquidity pools, highlighting analogies and differences between the two blockchains. We discover that about 60% of tokens are active for less than one day. Moreover, we find that 1% of addresses create an anomalous number of tokens (between 20% and 25%). We discover that these tokens are used as disposable tokens to perform a particular type of rug pull, which we call 1-day rug pull. We quantify the presence of this operation on both blockchains discovering its prevalence on the BNB Smart Chain. We estimate that 1-day rug pulls generated \$240 million in profits. Finally, we present sniper bots, a new kind of trader bot involved in these activities, and we detect their presence and quantify their activity in the rug pull operations.
If you cite this paper, please use the USENIX Security reference: Federico Cernera, Massimo La Morgia, Alessandro Mei, and Francesco Sassi. 2023. Token Spammers, Rug Pulls, and Sniper Bots: An Analysis of the Ecosystem of Tokens in Ethereum and in the Binance Smart Chain (BNB). In 32nd USENIX Security Symposium (USENIX Security 23). USENIX Association, Anaheim, CA, 3349–3366.
https://www.usenix.org/conference/usenixsecurity23/presentation/cernera
@@ 1 Introduction
The cryptocurrency market is loosely regulated [39, 4]. Even if policymakers are moving towards building a safer environment for cryptocurrency investors [56], it is a complex task, and needs time. Meanwhile, blockchain-related technologies evolve fast, and with the birth of the DeFi [67] investors begin to move from centralized exchanges (CEX) like Binance to decentralized exchanges (DEX). DEXes are distributed Applications (dApp) for trading that run on-chain powered by smart contracts. While regulating the standard cryptocurrency market is difficult, ruling the on-chain trading platform is even more challenging. Indeed, even if the web interface of a DEX can be shut down [3], its smart contracts are still reachable and working on the blockchain.
@@
DEX and DeFi dApp were born in the Ethereum blockchain, but DeFi services rapidly pop up on all the blockchains that support smart contracts. Although Ethereum plays a leading role in the DeFi world, with over \$68 billion locked in its smart contracts, the BNB Smart Chain or BSC (former Binance Smart Chain) proposes itself as a faster and cheaper alternative.
@@
Uniswap and PancakeSwap are the two most popular DEXes on Ethereum and BSC. They rely on the Automated Market Maker (AMM) model to handle the trading system. At the basis of the AMM model, there is the concept of liquidity pools, a smart contract that handles two tokens (trading pair) that the user can swap. Unlike CEX, where the platform defines the trading pairs, users can create their pair on DEXes and let the other users use it. However, as we will see in the following, some users abuse this freedom to carry out a series of malicious operations.
In this work, we conduct a longitudinal investigation of tokens and liquidity pools in the Ethereum and BSC blockchains. We start by parsing over 3 billion transactions of both blockchains, finding more than 1.3 million tokens and 1 million liquidity pools (Sec. 4). Then, we reconstruct their lifetime—the time from their creation to their last transfer, discovering that approximately 60% of the tokens have a lifetime shorter than one day (Sec. 5). Hence, we define them as 1-day tokens. A tiny fraction of addresses, just 1%, is responsible for creating more than 20% of the tokens (Sec. 6). Surprisingly, we also find that the tokens with a very short lifetime are actively traded on liquidity pools. Albeit this phenomenon is present on both blockchains, it is more widespread on BSC. Diving into this subset of tokens, we observe that a large fraction of liquidity pools used to trade the 1-day tokens show a malicious pattern that we call 1-day rug pull (Sec. 7). We analyze all the liquidity pools looking for this pattern, and we find 272,349 potential rug pulls on BSC and 21,742 on Ethereum. We estimate the cost of the operation and the gains of the organizers, finding that they earned approximately \$240 million with such activity (Sec. 7.2.1). Here, we see that the success rate of the 1-day rug pull is not very high (between 40% and 60%). However, given the simplicity and the very low cost of the operation, attackers can serially arrange the rug pulls and cover a series of unsuccessful operations with a single successful one. Finally, we study how this kind of operation evolved over time, discovering that the BSC has gradually surpassed Ethereum in terms of the number of operations and gains. Moreover, we find that the operations are more prevalent during two specific events: the 2020 Defi Summer and the 2021 Altcoin season (Sec. 7.2.2).
@@
  • Short lifetime tokens and Token spammers: We estimate the lifetime of the tokens on both blockchains. Discovering that about 60% of tokens last less than one day. Analyzing who creates the tokens, we observe that just 1% of addresses create an abnormal number of tokens (about 20-25% of tokens of the blockchains).
  • 1-day rug pulls: We investigate the presence of the rug pull pattern in 1-day tokens. We discover that on BSC, 81.2% of 1-day tokens listed on PancakeSwap have this pattern. We estimate the gains of the attackers, observing that even if the operation is very simple to arrange, given its cheap cost, it is profitable when performed serially.
  • The sniper bot 2.0: We find the presence of sniper bots (Sec. 8), a particular kind of trader bot that observes the blockchain’s mempool to buy newly listed tokens. To the best of our knowledge, we are the first to illustrate how this kind of trading bot works, detect their presence, and quantify their activity in the rug pull operations.
@@ 2 Ethereum and BNB Smart Chain
Ethereum [11] is a proof-of-work [^1] Switched to proof-of-stake from 2022-09-15 blockchain. Its native coin is the Ether (ETH), the second most popular cryptocurrency after Bitcoin (BTC), with a market cap of more than 210 billion USD. A key feature of Ethereum is smart contracts, pieces of code that execute in a decentralized way on-chain. Through smart contracts, it is possible to create new digital assets like (fungible) tokens and NFTs (non-fungible tokens).
@@
The tokens. Tokens are cryptocurrencies that can be exchanged or traded. They are created on top of the blockchain, and their mechanisms are defined using smart contracts. On Ethereum, the ERC-20 [26] standard defines the main properties of tokens. An ERC-20 compliant smart contract must implement a set of functions and events specified in the standard. These functions are reported in Table 5. Some of them are optional, in particular the name(), the symbol(), and the decimal() functions. On Ethereum, tokens and digital assets are held into accounts.
@@
Ethereum accounts. There are two kinds of accounts on Ethereum: Externally owned accounts (EOA) and contract accounts. EOAs consist of a pair of public and private keys generated with the Elliptic Curve Digital Signature Algorithm (ECDSA) [35]. An account is represented by its public address, a 42-character hexadecimal string obtained concatenating "0x" to the last 20 bytes of the Keccak-256 [23] hash of the public key. Instead, a contract account is tied to a smart contract. It is represented with an address in the same format as an EOA. A contract account is generated when a smart contract is deployed to the Ethereum blockchain.
@@
Transactions and fee. A transaction is an action that updates the whole Ethereum network. It can be used to move digital assets, deploy a smart contract, or invoke a smart contract. Executing a transaction has a cost, commonly called transaction fee. The fee is variable and depends on two main factors: The state of the network (if the network is heavily loaded, the fee is usually higher) and the complexity of the operation that the transaction triggers.
@@
Smart contract deployment. Smart contracts are programs that run on the Ethereum blockchain. They are written in a high-level programming language (e.g., Solidity [21]) and compiled into bytecode that runs on the Ethereum Virtual Machine (EVM) [25]. A smart contract can be deployed by sending a contract creation transaction from an EOA to the zero address2220x0000000000000000000000000000000000000000. The contract creation transaction contains the bytecode of the smart contract. A smart contract can also create new smart contracts. Since a smart contract can start a transaction only in response to a transaction that triggers it, an EOA must trigger the generation of a new smart contract.
Events and logs. To facilitate the tracking of internal states of smart contracts, Ethereum provides Events and an internal Logs register. Each time an action changes the internal state of a smart contract, it can fire an Event that notifies the change.
@@
EVM and EVM compliance. Ethereum is a distributed state machine that changes its state at each new block accordingly to a predefined set of rules. The EVM [65] is the entity that computes these changes in states. Other than Ethereum, other blockchains rely on the EVM (e.g., BNB Smart Chain [7], Avalanche [53], Fantom [27]), and they use one of the standard EVM or a custom one. These blockchains are called EVM-compliant. They run the same (or with minimal change) smart contract written for Ethereum, use the same convention for the address, and handle states the same way as Ethereum.
@@
The BNB Smart Chain. The BNB Smart Chain [7] (previously Binance Smart Chain) or BSC is a blockchain that was born in 2020. Its consensus is based on the PoSA [9] (Proof of Stake and Authority). The coin of both chains is the BNB (Build and Build, previously Binance Coin)—the third coin by market cap with over 46 billion of capitalization. As Ether for Ethereum, the BNB coin fuels the transactions on the BNB chain. Because of EVM compatibility, it is possible to create tokens in BSC similarly to Ethereum. However, in this case, tokens follow the BEP-20 standard instead of the ERC-20.
3 AMMs, Uniswap and its forks
Uniswap [1] is one of the most popular decentralized applications (dApp). According to DefiLlama [41], a popular DeFi statistics aggregator, Uniswap is the $5^{th}$ dApp by TVL (Total Value Locked, amount of money locked into smart contracts) with over 6 billion USD.
@@ 3 AMMs, Uniswap and its forks
Uniswap is the first dApp to use the AMM model successfully. This model relies on a mathematical formula to fix the price of assets and on the concept of liquidity pools and providers. A liquidity pool is a smart contract that contains two or more cryptocurrencies that the user can swap for another. Instead, a liquidity provider is a user who invests in the liquidity pool, providing cryptocurrencies to the smart contract. When a liquidity provider injects liquidity into the liquidity pool, the smart contract mints LP-tokens and gives them to the liquidity provider. The LP-token represents the share of the liquidity pool owned by the investor. When the liquidity provider desires to get back his cryptocurrencies, he transfers the LP-tokens to the smart contract. The latter burns the LP-tokens and provides the cryptocurrencies back to the investor. In Uniswap V2, each pool consists of a pair of ERC-20 tokens. The liquidity pool is divided into two parts, each containing a single token, and both have an equivalent value. Let a pool consist of $x$ token $A$ and $y$ token $B$. At each swap, the pool preserves $x*y$. When a user swaps $a$ token $A$ for token $B$ (the user adds token $A$ to the pool and takes token $B$ from the pool), $x$ increases by $a$ and $y$ decreases by $b$, where $b$ is computed so that $x*y$ does not change. The rate $a/b$ of the exchange depends on the ratio of $x$ and $y$ in the pool. Consequently, the swap operation changes the current exchange rate. The value of token $A$ decreases while the value of token $B$ increases, and the two parts maintain the same value.
Because of the success of this model, the popularity of Uniswap and its open-source smart contracts [62], more than 50 protocols were born on several blockchains by forking Uniswap smart contracts in the last years. Uniswap is on its third version, but all its forks belong to the second version since the third one is under a Business Source License. For this reason, in this work, we focus on Uniswap V2 and its forks. One of the most popular forks of Uniswap is PancakeSwap, which lives on BSC. It is the $1^{th}$ dApp by TVL on this blockchain with over 4 billion USD locked in its smart contracts.
@@ 4 The Datasets
For our investigation, we build two different datasets: The Token Dataset, which contains all the ERC-20 (resp. BEP-20) tokens created, and the Liquidity Pool Dataset, which contains data about liquidity pools. Each dataset has two versions, one with data from the Ethereum blockchain and the other from the BNB Smart Chain.
We consider the whole history of both blockchains from their inception to March 2022. For the Ethereum blockchain, we process all the blocks from block 0 (2015-07-30) to block 14340000 (2022-03-07). For the BSC blockchain from block 0 (2020-04-20) to block 15854000 (2022-03-07). Given the large amount of data and the need to parse the entire blockchains multiple times, for performance reasons and to avoid overloading public nodes (e.g., nodes provided by Binance [6] and Infura [34]) or services (e.g., BscScan or Etherscan), we host and run an Ethereum and a BNB Smart Chain node. Finally, to query the blockchains and process the data, we use the Web3 [46] and the Ethereum-etl [44] Python libraries. Web3 is a collection of libraries that allow the interaction with a local or remote EVM-compliant node. Ethereum-etl allows extracting information from EVM-compliant blockchains and exporting it into formats like CSV or JSON. The data collection phase was performed on an Ubuntu 20.04 machine with AMD EPYC 7301 (16-Core Processor, 2.80 GHz), 1 TB of RAM, and 4 TB SATA SSD with 560/530 MB/s read and write speed. Data processing took between 24 and 72 hours each time we parsed the entire blockchain, depending on the kind of data retrieved.